# The radiative decay of scalar glueball from lattice QCD

### v2

@@ V SUMMARY
We first revisit the radiative $J/\psi$ decay to the scalar glueball and derive the electromagnetic form factor $E_{1}(0)=0.0618(49)~{}\text{GeV}$ in the continuum limit, which gives the partial decay width $\Gamma(J/\psi\to\gamma G)=0.449(44)~{}\text{keV}$ and the branching fraction $\mathrm{Br}(J/\psi\to\gamma G)=4.8(5)\times 10^{-3)}$ when $m_{G}\approx 1.635(62)~{}\text{GeV}$ and the total width $\Gamma=92.6(1.7)~{}\text{keV}$ of $J/\psi$ are used. These results confirm the previous lattice results $0.35(8)~{}\text{keV}$ and $3.8(9)\times 10^{-3}$ [28] and support $f_{0}(1710)$ to be a candidate for the scalar glueball or to have a large component of it.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12175063, No. 12175073, No. 12222503, No. 11935017, No. 12293060, No. 12293065, No. 12070131001 (CRC 110 by DFG and NSFC)). JL is also supported by the Natural Science Foundation of Basic and Applied Basic Research of Guangdong Province under Grants No. 2023A1515012712. CY also acknowledges the support by the National Key Research and Development Program of China (No. 2020YFA0406400) and the Strategic Priority Research Program of Chinese Academy of Sciences (No. XDB34030302). LG also acknowledges the support by the Hunan Provincial Natural Science Foundation (No. 2023JJ30380). WQ also acknowledges the support by the Hunan Provincial Natural Science Foundation (No. 2024JJ6300) and the Scientific Research Fund of Hunan Provincial Education Department (No. 22B0044). The numerical calculations are carried out on the GPU cluster at Hunan Normal University. Our matrix inversion code is based on QUDA libraries [56] and the fitting code is based on lsqfit [57].

### v3

@@ V SUMMARY
We first revisit the radiative $J/\psi$ decay to the scalar glueball and derive the electromagnetic form factor $E_{1}(0)=0.0677(48)~{}\text{GeV}$ in the continuum limit, which gives the partial decay width $\Gamma(J/\psi\to\gamma G)=0.578(86)~{}\text{keV}$ and the branching fraction $\mathrm{Br}(J/\psi\to\gamma G)=6.2(9)\times 10^{-3}$ when $m_{G}\approx 1.635(62)~{}\text{GeV}$ and the total width $\Gamma=92.6(1.7)~{}\text{keV}$ of $J/\psi$ are used. Comparing to the previous lattice results $0.35(8)~{}\text{keV}$ and $3.8(9)\times 10^{-3}$ [30], this work includes the data at finer lattice spacing and uses wall source propagators, which effectively improve the signal. This provides more solid support for $f_{0}(1710)$ being a candidate for the scalar glueball or having a large component of it.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12175063, No. 12175073, No. 12222503, No. 11935017, No. 12293060, No. 12293062, No. 12293065, No. 12070131001 (CRC 110 by DFG and NSFC)). JL is also supported by the Natural Science Foundation of Basic and Applied Basic Research of Guangdong Province under Grants No. 2023A1515012712. CY also acknowledges the support by the National Key Research and Development Program of China (No. 2020YFA0406400) and the Strategic Priority Research Program of Chinese Academy of Sciences (No. XDB34030302). LG also acknowledges the support by the Hunan Provincial Natural Science Foundation (No. 2023JJ30380). WQ also acknowledges the support by the Hunan Provincial Natural Science Foundation (No. 2024JJ6300) and the Scientific Research Fund of Hunan Provincial Education Department (No. 22B0044). The numerical calculations are carried out on the GPU cluster at Hunan Normal University. Our matrix inversion code is based on QUDA libraries [60] and the fitting code is based on lsqfit [61].